# Interpreting the Spectro-Temporal Properties of the Black Hole Candidate Swift J151857.0-572147 during its First Outburst in 2024

### v1

@@ 4 Discussions
The Galactic black hole Swift J151857.0-572147 started an outburst recently in March 2024. We have used Insight-HXMT data for our both timing and spectral studies from 2024 March 04 to 2024 March 17. Using the $0.01$ sec time-binned light curves from the three instruments of HXMT (LE, ME, and HE), we studied the source’s timing properties. We also searched for the energy dependence of LFQPOs by producing light curves in 7 different energy ranges within the HE band. Along with these, we searched for HFQPOs from all the light curves from LE, ME, and HE by making $1$ ms time-binned light curves. We then examined the combined LE+ME+HE spectra in the $2-100$ keV broad energy band to learn more about the spectral characteristics of this source using the spectra files from these three instruments.
@@
For stellar-mass black holes, quasi-periodic oscillation is one of the most significant and frequent occurrences. We examined 186 exposures in total for this recently found source (62 for each of LE, ME, and HE). Nevertheless, incorrect light curve production occurred in 2 LE exposures. A total of 184 light curves for LE, ME, and HE were obtained. The details are listed in Table 2. We discovered that QPO was not present in each of these light curves. The details on QPO properties are listed in Table 3. Over this brief analysis period, the QPO frequency has rapidly changed. Even in a single day, there was an increase in the QPO frequency ($\nu_{qpo}$). The results section contains a general discussion of the QPO frequency’s evolution. Type-C QPO nature is identified from the fluctuation of QPO frequency, (%)RMS, and $Q$-factor. One thing we would like to discuss here that the difference in RMS value (in Table 3) for LE, ME, and HE is due to the large variation in background counts in these three bands. As we can notice in Table 2, the background count in the HE band is almost equivalent to the source count in the HE band for this source. We recently studied the outburst properties of the BHC Swift J1727.8-1613 using HXMT data (Chatterjee et al. 2024). However, there was a great difference between the source and background count rates in all three bands. This is the case here for the LE and ME bands. For both sources, the same reduction method is used.
@@
Even though the QPOs have been thoroughly examined in the literature, further modeling is necessary to understand their origin. Here, we want to concentrate on the physical scenario that explains how shock instabilities in advective flows near black holes (BHs) give rise to QPOs. According to Chakrabarti (1989), accretion onto BH is a transonic flow with the potential for numerous sonic locations. The companion’s matter doesn’t need to be exclusively Keplerian. A supply of matter with an angular momentum distribution that differs from the Keplerian one may exist. This is the sub-Keplerian component. Infalling matter with a smaller angular momentum component accretes over a free-fall timescale. At a certain distance from the black hole, this matter might nearly stop due to the counterbalance between the centrifugal force and gravitational force, undergo a shock transition, and form a post-shock region. Standing shocks form according to the flow properties (Chakrabarti 1989; Singh et al. 2022 and references therein). The spectral and temporal properties of BHs that have been seen can be effectively described by this method (Debnath et al. 2014; Mondal et al. 2014, Chatterjee et al. 2020, 2021, 2023). This shock may not be stable at the outer edge over time. There could be oscillations in the CENBOL boundary, which can be caused due to either of two reasons:
@@
(i) According to Chakrabarti (1989), the satisfaction of the Rankine-Hugoniot condition makes the boundary of the shock stable and steady. However, if this condition is not satisfied, the shock could oscillate at the outer boundary to find local equilibrium. This could produce variabilities in the light curves.
@@
(ii) Molteni et al. (1996) state that the presence of cooling may cause the shock to oscillate. QPOs emerge during the oscillation when the heating timescale and the cooling timescale caused by the Comptonization process match (see Chakrabarti et al. 2015).
The QPO frequency is given by,
$\nu_{qpo}=\frac{c^{3}}{2GM_{BH}}\frac{1}{RX_{s}(X_{s}-1)^{1/2}}Hz,$ (1)1( 1 )
where the following are represented, respectively: $c$, $G$, $M_{BH}$, $X_{s}$, and $R$; these are the gravitational constant, the speed of light, the mass of the BH, the shock location, and the ratio of matter densities in post-shock to pre-shock regions ($\rho_{+}/\rho_{-}$).
Figure 14: Variation of shock properties with time in accordance with $\nu_{qpo}$. In (a) the variation of $\nu_{qpo}$ with time and in (b) we show the variations of the shock location ($X_{s}$) with time.
@@
We have already retrieved the QPO frequency ($\nu_{qpo}$) from our timing analysis. We calculated the shock location during the outburst using the relation mentioned above. Chakrabarti et al. (2005) state that depending on the flow parameters causing shocks, $X_{s}$ can be anywhere over $10~{}r_{s}$. When the spectral nature of an outburst is hard, the shock forms far away $\sim 1000~{}r_{s}$, and it gradually becomes small in the following days as cooling increases (Mondal et al. 2015). Based on Eq. 1, we discovered that during the beginning of the outburst, the shock was located at a distance of $\sim 100~{}r_{s}$ from the BH (see Figure 14b). Then, as $\nu_{qpo}$ increased, the shock moved inwards, suggesting cooling was in progress. After a few days, the shock became stable. Table 3 provides the values for the shock location (columns 11–13).
@@
The light curve from MAXI/GSC was absent due to the proximity of another source in Cir X-1. Thus, from the HXMT extracted light curves and HR variations, we can roughly say that the source transitioned past its hard state at the start of our analysis period. It was at the end of its harder state (HS and HIMS). Our spectral analysis result also confirms this. From the variation of the photon index, we can say that the source was already in the end phase of its HIMS state and was transitioning towards its SIMS state. Our estimated shock location also agrees with this designation. At the start of our analysis period, shock was at a distance, which suggests the source has already transitioned past its hard state and may be at the end of its hard intermediate state. Then the shock moved inward, suggesting now the source is making a transition towards the softer states via SIMS and SS. On several days, we did not notice any QPO signature from the start of our analysis to the end in any of the three bands. This could be because the infall and cooling timescales did not match for the shock. However, this could be also due to the data. For several days, QPO is only present in one of the three bands, whereas on some days it is present in two bands. We find that on the last day of our analysis period, the $\Gamma$ became quite high, which suggests that the source transitioned into the soft state (SS) that day. Thus, we did not find QPO in any band on that day. The absence of QPOs and high $\Gamma$ confirms the SS as inferred in Mondal et al. (2024) using IXPE and NuSTAR observations of the source.
We also found that high-frequency QPO was absent during the entire analysis duration of the outburst. Although the HFQPO phenomenon is not very common, its absence could be because the disk did not proceed very close to the compact object to produce variabilities with high frequency. Or, it could be due to the reason that the photon detection by HXMT instruments is not sufficient for this phenomenon. We also checked the energy dependence of LFQPOs in the HE band for those exposures in which LFQPO was present in the full energy band. The energy dependence of QPOs could give valuable insight into the origin of the QPO. In Ma et al. (2023) paper, they reported that LFQPO was present till very high energy, which suggested that the origin of the QPO could be from the precession of the jet. Examining all of those 31 exposures, we find that LFQPO was present till $48$ keV, above which there is no prominent or weak QPO nature, either. In the $27-35$ keV band, the nature of LFQPOs was stronger than in the $35-48$ keV band. The shock was quite strong in the case of the BHC Swift J1727.8-1613 (Chatterjee et al. 2024), which could produce variabilities up to higher energies. However, for this source, the shock was not very big, which suggests that it has already cooled down by the process of inverse-Comptonization. Thus, it could only produce variabilities in hard photons up to $48$ keV. This also supports our claim about the origin of the quasi-periodic oscillations, observed in the light curves.
@@
Unlike the BHC Swift J1727.8-1613, this source has shown a very high value of hydrogen column density ($N_{H}$), using both combinations of models. $N_{H}$ varied in the range of $(4.3-6.5)\times 10^{22}$ cm-2 and $(5.1-6.3)\times 10^{22}$ cm-2 for Model-1 and Model-2, respectively. This is very high, considering Galactic black holes. This indicates some absorption local to the source, which could be due to the outflows from the disk or the presence of some blobs along the line of sight (see Neilsen & Homan, 2012; Mondal & Jithesh, 2023). To confirm this, we need a detailed study of the outflow/jet properties of the source.
@@ 5 Summary and Conclusions
We have studied the timing and spectral properties of the very first outburst of BHC Swift J1727.8-1613 in 2024. Using Insight-HXMT LE, ME, and HE exposure average light curve data, we present the evolution of the light curve and its hardness ratio across our full analysis period. For our investigation, we selected the 7 observation IDs using the Insight-HXMT data, publicly available during the analysis. For timing analysis, we employed all of the exposures from those observation IDs, and for spectrum analysis, we employed selective exposures, respectively. We produced a power density spectrum and used $0.01~{}s$ time-binned light curves from the three HXMT instruments, i.e., LE, ME, and HE, to study the QPO properties. We used the Lorentzian model to obtain the QPO properties. We also studied energy-dependent QPO by producing HE light curves in seven different energy bands. We extracted the energy-dependent QPO properties in the same way we did for the LE, ME, and HE light curves in the full band. Apart from these, we also produced $0.001$ s time-binned light curve to search for high-frequency QPOs. We employ LE + ME + HE spectrum files in the broad $2-100$ keV energy band for spectral analysis. We found that the models i) constant*tbabs*(diskbb + broken power-law) and ii) constant*tbabs*(diskbb + pexrav) fit the spectra for the best statistics. Based on our investigation, we cdiskonclude that:
@@
v) LFQPOs were present up to 48 keV, above which we did not find the presence of LFQPO for all the exposures.
@@
viii) The hydrogen column density varied in the range of $N_{H}\sim(4.3-6.9)\times 10^{22}$ cm-2 in accord with the estimation by Mondal et al. (2024). This could be due to the presence of outflows from the disk or some blobs along the line of sight.
@@ 7 Acknowledgements
We thank Dr. Lian Tao of the Institute of High Energy Physics (IHEP), Chinese Academy of Sciences (CAS) for publicizing the Insight-HXMT data by request. We also thank Prof. Mutsumi Sugizaki of the National Astronomical Observatory, Chinese Academy of Sciences (NAOC) for providing fruitful information about the absence of MAXI/GSC daily average light curve on the source. We sincerely thank Prof. John A. Tomsick of the Space Sciences Lab, University of California, Berkeley, USA for providing suggestions in the draft.
KC acknowledges support from the SWIFAR postdoctoral fund of Yunnan University. SPS and SM acknowledge the Ramanujan Fellowship (# RJF/2020/000113) by SERB-DST, Govt. of India for this research. CBS is supported by the National Natural Science Foundation of China under grant no. 12073021.

### v2

@@ 4 Discussions
The Galactic black hole Swift J151857.0-572147 started an outburst recently in March 2024. We have used Insight-HXMT data for our both timing and spectral studies from 2024 March 04 to 2024 March 17. Using the $0.01$ sec time-binned light curves from the three instruments of HXMT (LE, ME, and HE), we studied the source’s timing properties. We also searched for the energy dependence of LFQPOs by producing light curves in 7 different energy ranges within the HE band. Along with these, we searched for HFQPOs in all the PDS from LE, ME, and HE using $1$ ms time-binned light curves. We then examined the combined LE+ME+HE spectra in the $2-100$ keV broad energy band to learn more about the spectral characteristics of this source using the spectra files from these three instruments.
@@
For stellar-mass black holes, quasi-periodic oscillation is one of the most significant and frequent occurrences. We examined 186 exposures in total for this recently found source (62 for each of LE, ME, and HE). Nevertheless, incorrect light curve production occurred in 2 LE exposures. A total of 184 light curves for LE, ME, and HE were obtained. The details are listed in Table 2. We discovered that QPO was not present in each of these light curves. The details on QPO properties are listed in Table 3. Over the analysis period of $\sim 13$ days, the QPO frequency has rapidly changed. In our analysis period, the $\nu_{qpo}$ showed variation approximately from 3.3 to 7.4, 3.2 to 9, and 3.2 to 7 Hz in LE, ME, and HE bands, respectively. Even in a single day, there was a change in the QPO frequency ($\nu_{qpo}$). The results section contains a general discussion on the evolution of QPO frequencies. Type-C QPO nature is identified from the fluctuation of QPO frequency, (%)RMS, and $Q$-factor. One thing we would like to discuss here that the difference in RMS value (in Table 3) for LE, ME, and HE is due to the large variation in background counts in these three bands. As we can notice in Table 2, the background count in the HE band is almost equivalent to the source count in the HE band for this source. The high background count rate in the HE band could be due to the contribution of the close proximity of the other source Cir X-1 and the high effective area of the detector in this band. However, it was previously reported by Peng et al. (2024) that Cir X-1 was mainly present in the soft spectral state during the outburst of Swift J151857.0-572147. Thus, it could only have contributed to the LE band. Thereby the reason for this high background rate can not be firmly concluded.
@@
Even though the QPOs have been thoroughly examined in the literature using observations from other sources, further modeling is necessary to understand their origin and connection with the spectral properties. Here, we want to concentrate on the physical scenario that explains how shock instabilities in advective flows near black holes (BHs) give rise to QPOs. It is already explained in the introduction how the shock oscillations in the TCAF model explains the origin of the QPOs. This shock may not be stable at the outer edge over time. There could be oscillations in the CENBOL boundary, which can be caused due to either of two reasons:
@@
(i) According to Chakrabarti (1989), the satisfaction of the Rankine-Hugoniot condition makes the boundary of the shock stable and steady. However, if this condition is not satisfied (Ryu et al. 1997), the shock could oscillate at the outer boundary. It could produce variabilities in the light curves.
@@
(ii) Molteni et al. (1996) stated that the presence of cooling may cause the shock to oscillate. QPOs emerge during the oscillation when the compressional heating timescale and the cooling timescale due to inverse Comptonization process match (see Chakrabarti et al. 2015).
@@
Depending on the flow parameters causing shocks, $X_{s}$ can be anywhere over $10~{}r_{s}$. When the spectral nature of an outburst is hard, the shock forms far away $\sim 1000~{}r_{s}$, and it gradually becomes small in the following days as cooling increases (Mondal et al. 2015). For this outburst, the shock was far from the BH at the start of our analysis period. As the spectral nature of the outburst was softening, the shock moved inwards, suggesting cooling was in progress. As the shock moved inwards, the QPO frequency increased.
@@
Although MAXI/GSC observed the source, it was not identified as a new source due to the proximity of another source, called Cir X-1. Thus, from the HXMT extracted light curves and HR variations, we may say that the source transitioned past its HS at the start of our analysis period and was already moved to the intermediate state. The variation of the photon index supports the above spectral state. At the start of our analysis period, the shock was at a distance, which suggests that the source had already completed its HS and was in the intermediate state. Later, the shock moved inward, suggesting the source is making a transition towards the soft state. The non-identification of type-B QPO does not help to designate a transition between the HIMS and SIMS, and thus we consider this overall observation period belonging to the intermediate state. On several exposure Ids, we did not notice any QPO signature from the start of our analysis to the end in any of the three bands (Table 2 & 3 for correspondence). This could be because of the mismatch of the heating and cooling timescales at the shock. For several exposure Ids, QPO is only present in one of the three bands, whereas, for some of them it was present in two bands (Table 2 & 3 for correspondence). We find that on the last day of our analysis period, the $\Gamma$ became high ($>3.3$ from all the three models). These values suggest that the source transitioned into the SS that day (Remillard & McClintock 2006). Thus, we did not find QPO in any band on that day. The absence of QPOs and high $\Gamma$ agrees with the SS as inferred in Mondal et al. (2024) using joint IXPE and NuSTAR observations of the source. As mentioned previously in the result section, we have not noticed typical variations of the $\Gamma$, $Norm_{diskbb}$, $T_{in}$ and we have also found a broad range of the spin and inclination from the spectral analysis. We speculate that this is due to the contribution from the Cir X-1 source, albeit it was in the low energy band.
We did not find the presence of any high-frequency QPO during the entire analysis period of the outburst. The HFQPO phenomenon is not very common. To date, HFQPO has been observed in a few sources only, e.g., GRO J1655-40 (Remillard et al. 1999; Strohmayer 2001b; Remillard et al. 2006), H1743-322 (Homan et al. 2005; Remillard et al. 2006), XTE J1550-564 (Homan et al. 2001; Miller et al. 2001; Remillard et al. 2002a), and GRS 1915 + 105 (Morgan, Remillard & Greiner 1997; Strohmayer 2001b; Belloni et al. 2006). This suggests that this is not a very common phenomenon, like LFQPO in BHXRBs. Its absence could be because the disk did not proceed very close to the compact object to produce variabilities with high frequency. Detection of HFQPOs requires high timing resolution, large effective area, both of which are synonymous with the Insight-HXMT satellite. However, the photon statistics should be very strong to detect HFQPOs that requires very high signal-to-noise (S/N) ratio, especially in soft energy band, as it is generally thought to be produced by the disk when it is very close to the black hole. Considering this source was at proximity of another source, this may not have satisfied. These could be plausible causes for the non-detection of HFQPOs for this outburst. We also studied the energy dependence of LFQPOs in the HE band for those exposures in which LFQPO was present in the full energy band. The energy dependence of QPOs could give valuable insight into the origin of the QPO. Ma et al. (2023) reported that LFQPO was present till very high energy, which suggested that the origin of the QPO could be from the precession of the jet. Examining all 31 exposures, we find that LFQPO was present till $48$ keV, above which there is no prominent or weak QPO nature, either. In the $27-35$  keV band, the nature of LFQPOs was stronger than in the $35-48$ keV band. Such observational findings can be explained from TCAF model scenario, where in the intermediate states CENBOL shrinks, due to the increase in cooling effects (see, Mondal et al. 2015), therefore the spectral break energy permissible for inverse-Compton scattering also decreases, which is the case for the present source. Thereby, much higher energy photons possibly could not contribute to the observed QPOs. There could be another explanation for the absence of QPOs at high energies. We have noticed that the background count rates are higher for the HE band and are very comparable to the source count. Due to the high background, there could be less contribution from the source at higher energy bands. This could also explain why we do not observe QPOs at these HE bands. Since the source is already in the SIMS, the RMS amplitude is generally lower than in the HS. Therefore, the upper limit on the RMS amplitude can not be constrained from the present data sets, requiring a detailed study with the complete cycle of an outburst.
@@
The source Swift J151857.0-572147 has shown a very high value of $N_{H}$, using all three combinations of models. The $N_{H}$ varied in the range $(4.3-6.5)\times 10^{22}$  cm-2, $(5.1-6.3)\times 10^{22}$ cm-2, and $(5-6.9)\times 10^{22}$ cm-2 for Model-1, Model-2, and Model-3, respectively. The average column density was $\sim 5.6\times 10^{22}~{}cm^{-2}$. This value is significantly high, compared to other Galactic black holes. For example, some BHCs Swift J1727.8-1613, MAXI J1803-298, GX 339-4, and Swift J1357.2-0933 the $N_{H}$ varied in the range $(0.1-0.5)\times 10^{22}~{}cm^{-2}$ (Liu et al. 2024; Debnath et al. 2024; Chatterjee et al. 2024), $(0.2-0.5)\times 10^{22}~{}cm^{-2}$ (Adegoke et al. 2024; Jana et al. 2022), $0.5\times 10^{22}~{}cm^{-2}$ (Motta et al. 2009), and $0.13\times 10^{22}~{}cm^{-2}$ (Mondal & Chakrabarti 2019). For these sources, there was no local absorption and thus the values were close to the Galactic hydrogen column density (HI4PI Collaboration 2016). This has also been observed for other BHXRBs for which there was no local absorption to the source. This indicates some absorption local to the source, which could be due to the outflows from the disk or the presence of some blobs along the line of sight (see Neilsen & Homan, 2012; Mondal & Jithesh, 2023). To confirm this, we need a detailed study of the outflow/jet properties of the source.
@@ 5 Summary and Conclusions
We have studied the timing and spectral properties of the very first outburst of the BHC Swift J151857.0-572147 in 2024. Using Insight-HXMT LE, ME, and HE exposure average light curve data, we present the evolution of the light curve and its hardness ratio across our full analysis period from March 04 2024 (MJD 60373) to March 17 2024 (MJD 60386). For our investigation, we selected the 7 observation IDs using the Insight-HXMT data, publicly available during the analysis. For timing analysis, we employed all of the exposures from those observation IDs, and for spectrum analysis, we employed selective exposures, respectively. We produced a power density spectrum and used $0.01~{}s$ time-binned light curves from the three HXMT instruments, i.e., LE, ME, and HE, to study the QPO properties. We used the Lorentzian model to obtain the QPO properties. We also studied energy-dependent QPO by producing HE light curves in seven different energy bands. We extracted the energy-dependent QPO properties in the same way we did for the LE, ME, and HE light curves in the full band. Apart from these, we also produced $0.001$ s time-binned light curve to search for high-frequency QPOs. We employ LE + ME + HE spectrum files in the broad $2-100$ keV energy band for spectral analysis. We found that the models i) constant*tbabs*(diskbb + broken power-law) and ii) constant*tbabs*(diskbb + pexrav) fit the spectra for the best statistics. Based on our investigation, we conclude that:
@@
v) LFQPOs were present up to 48 keV, above which we did not find the presence of LFQPO for any of the exposures.
@@
viii) The average hydrogen column density was high with $N_{H}\sim 5.6\times 10^{22}$ cm-2 in accordance with the estimation by Mondal et al. (2024), Peng et al. (2024). This could be due to the presence of outflows from the disk or some blobs along the line of sight, which prompts a further detailed study.
@@ 7 Acknowledgements
We thank Dr. Lian Tao of the Institute of High Energy Physics (IHEP), Chinese Academy of Sciences (CAS) for publicizing the Insight-HXMT data by request. We also thank Dr. Mutsumi Sugizaki of the Kanazawa University, Japan for providing fruitful information about the absence of MAXI/GSC daily average light curve on the source. We sincerely thank Prof. John A. Tomsick of the Space Sciences Lab, University of California, Berkeley, USA for providing suggestions on the draft. CBS is supported by the National Natural Science Foundation of China under grant no. 12073021. SPS and SM acknowledge the Ramanujan Fellowship (RJF/2020/000113) by SERB/ANRF-DST, Govt. of India for this research. KC acknowledges support from the SWIFAR postdoctoral fund of Yunnan University.